OpenStudy (anonymous):

A master plumber can complete a job in 6 working days. After he has worked for 2 days, he was called away for a new job, and another plumber completed the job in 10 additional days. How long would it take the second plumber to do the job if he worked alone? Please show me the steps.

OpenStudy (anonymous):

First, we have to figure out how much of the job the second plumber completed. If the master plumber worked for 2 days, and would have completed it in 6 days, that means he completed \(\frac{2}{6}\) of the job, which means that the second plumber still had to complete \(\frac{4}{6}\) of the job. If the second plumber took 10 days to complete \(\frac{4}{6}\) of the job, how long would it take him to complete \(\frac{6}{6}\) of the job?

OpenStudy (anonymous):

thanks